# Random matrix model of Kolmogorov-Zakharov turbulence

### v1

@@ IV.3 KZT and Anderson localization
Here $\varepsilon_{n}$ are random on-site energies randomly and homogeneously distributed in the interval $W/2\leq\varepsilon_{n}\leq W/2$, the hopping takes place only on nearest sites with unit amplitude, $f$ describes a static Stark field. At $f=\beta=0$ this system represents the one-dimensional Anderson model with exponentially localized eigenmodes (see e.g. kramer ; mirlin ). The localization length at the middle of the energy band is $\ell=96/W^{2}$. At moderate nonlinearity with $\beta\sim 1$ and $f=0$ this Anderson localization is destroyed and an unbounded subdiffusive spreading of wave packet takes place dls2008 ; flach . This spreading is preserved in presence of a moderate Stark field garcia . In rmtprl it was shown that in such a system of finite size $N$ the dynamical thermalization takes place at a moderate nonlinearity $\beta\sim 1$ and a moderate Stark field $f$ with the steady-state thermal distribution corresponding to energy equipartition (2).
Figure 8: $\rho_{m}$ versus $E_{m}-E_{1}$ in a double logarithmic representation for the DANSE model with diagonal shift parameter $f=0.125$, disorder strength $W=4$, $N=256$, $\gamma=\sigma=0.01$, $\beta=1$, and pumping/absorption/initial condition as in Fig.  1. The values of $\rho_{m}$ have been obtained from the time average $\rho_{m}(t)=\langle|C_{m}(\tau)|^{2}\rangle$ for $t/2\leq\tau\leq t$ with $t=2^{22}$ (red ++), $t=2^{20}$ (green $\times$), $t=2^{18}$ (blue *) and $t=2^{16}$ (pink $\square$). The straight line shows the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the data for $t=2^{22}$ and the fit range $0.01(E_{N}-E_{1})\leq(E_{m}-E_{1})\leq 0.5(E_{N}-E_{1})$ with $A=1.00\pm 0.04$, $s_{0}=1.02\pm 0.02$ (and $E_{1}=-1.67,\,E_{N}=33.42,\,E_{N}-E_{1}=35.09$). The total norm is $\kappa=13.86$ (for $t=2^{16}$), $17.16$ (for $t=2^{18}$), $19.58$ (for $t=2^{20}$), $21.19$ (for $t=2^{22}$).
@@ V Discussion
We note that the dynamical thermalization has been recently observed in the experiments with multimode fibers (see e.g. fiber1 ; fiber2 ; fiber3 ; fiber4 ; fiber5 ) and we assume that the RMM system discussed here for the KZ turbulence can be realized in fibers with a cross-section of a chaotic 2D billiard. Such quantum billiards have many properties of quantum chaos bohigas ; haake being similar to the RMT case studied here. As discussed in rmtprl , we think that a most optimal billiard section of fiber is a D-shape one (a circle with cut) where the classical dynamics is know to be chaotic (see e.g. cao ).

### v2

@@ IV.3 KZT and Anderson localization
Here $\varepsilon_{n}$ are random on-site energies randomly and homogeneously distributed in the interval $-W/2\leq\varepsilon_{n}\leq W/2$, the hopping takes place only on nearest sites with unit amplitude, $f$ describes a static Stark field. At $f=\beta=0$ this system represents the one-dimensional Anderson model with exponentially localized eigenmodes (see e.g. kramer ; mirlin ). The localization length at the middle of the energy band is $\ell=96/W^{2}$. At moderate nonlinearity with $\beta\sim 1$ and $f=0$ this Anderson localization is destroyed and an unbounded subdiffusive spreading of wave packet takes place dls2008 ; flach . This spreading is preserved in presence of a moderate Stark field garcia2 . In fact the spreading is subdiffusive such that the second moment grows as $\langle n^{2}\rangle\propto t^{\nu}$ with the numerical exponent value $\nu\approx 0.3-0.4$ in 1D dls2008 ; flach ; garcia2 . This value is smaller than $\nu=0.5$ following from the random phase approximation basko ; nazarmdpi . In rmtprl it was shown that in such a system of finite size $N$ the dynamical thermalization takes place at a moderate nonlinearity $\beta\sim 1$ and a moderate Stark field $f$ with the steady-state thermal distribution corresponding to energy equipartition (2).
Figure 8: $\rho_{m}$ versus $E_{m}-E_{1}$ in a double logarithmic representation for the DANSE model with diagonal shift parameter $f=0.125$, disorder strength $W=4$, $N=256$, $\gamma=\sigma=0.01$, $\beta=1$, and pumping/absorption/initial condition as in Fig.  1. The values of $\rho_{m}$ have been obtained from the time average $\rho_{m}(t)=\langle|C_{m}(\tau)|^{2}\rangle$ for $t/2\leq\tau\leq t$ with $t=2^{22}$ (red $+$), $t=2^{20}$ (green $\times$), $t=2^{18}$ (blue *) and $t=2^{16}$ (pink $\square$). The straight line shows the power law fit $\rho_{m}=A\,(1+E_{m})^{-s_{0}}$ using the data for $t=2^{22}$ and the fit range $0.01(E_{N}-E_{1})\leq(E_{m}-E_{1})\leq 0.5(E_{N}-E_{1})$ with $A=1.00\pm 0.04$, $s_{0}=1.02\pm 0.02$ (and $E_{1}=-1.67,\,E_{N}=33.42,\,E_{N}-E_{1}=35.09$). The total norm is $\kappa=13.86$ (for $t=2^{16}$), $17.16$ (for $t=2^{18}$), $19.58$ (for $t=2^{20}$), $21.19$ (for $t=2^{22}$).
@@ V Discussion
We note that the dynamical thermalization has been recently observed in the experiments with multimode fibers (see e.g. fiber1 ; fiber2 ; fiber3 ; fiber4 ; fiber5 ) and we assume that the RMM system discussed here for the KZ turbulence can be realized in fibers with a cross-section of a chaotic 2D billiard. Such quantum billiards have many properties of quantum chaos bohigas ; haake being similar to the RMT case studied here. As discussed in rmtprl , we think that a most optimal billiard section of fiber is a D-shape one (a circle with cut) where the classical dynamics is know to be chaotic (see e.g. cao ). We expect that in multimode fibers the KZT regime can be realized by a continuous laser pumping of low energy modes while the light at high energy modes will escape from a fiber due to high angle collisions with fiber perimeter thus leading to a stationary KZT flow from low to high energies.